Booster: 1.3s cycle today, slip stacking next week 4+6

MITPC: running, tie in RWM from SciBath

ANNIE: slow progress

SciBath: Kerberos issues, Rex(says Kevin fixed it)

RWM: not digitizing?

MicroBooNE: DAQ uptime >97%, need to request no extra heating

MiniBooNE: leave temperatures alone

Toroids: Aisha re-did calibrations on toroids early last week T60 & T75. T60 ended up remaining the same. For T75, calibrations moved up by a few %. The two agree within a few percent (toroids: to 0.01 are the same). Check over all range.

Data loggers IFBEAM: Jim Patrick, restart loggers to get new constants.
Between data loggers and front end, posting at high resolution. Removes double counting/no counting, happens sometimes, but at correct level.

For the latched devices: some repeats. Questionable if we need them. Homework: understand how frontend orders requests.